OpenStudy (anonymous):

Need Help!!!!!!!!!! Will Give A Medal!!!!!!!!!! On scrap paper, determine if the triangle with the following vertices is a right triangle. A(-3, 5), B(1, 4), and C(-1, 0). Tell whether or not the triangle is a right triangle. Explain how you arrived at your answer.

OpenStudy (sidsiddhartha):

A=(-3,5) B=(1,4) so slope of line ab=m1=(y2-y1)/(x2-x1)=4-5/1+3=-1/4 slope of line BC=m2=0-4/-1-1=2 so m1*m2=-1/2 which is not equal to 1 so its not a right triangle